OpenStudy (deathslasher99):

Solve the following system of linear equations giving the solution as a coordinate pair in the form (x,y): y = 7x + 7 y = 8x - 3

OpenStudy (deathslasher99):

Can you tell me when you get the answer please, i just need to verify what i am doing wrong Thank you

OpenStudy (welshfella):

as they both have y = then we can equate the right hand sides so 7x + 7 = 8x - 3

OpenStudy (welshfella):

this is simple to solve first subtract 7x from both sides then add 3 to both sides and you'll get some value = x

OpenStudy (welshfella):

ill do the first part 7x + 7 - 7x = 8x - 7x - 3 7 = x - 3 x = ?

OpenStudy (deathslasher99):

I dont understand your process? Do you want me to do 7x + 7 + 3x=8x- 7x -3 + 3? Is this what you want me to do?

OpenStudy (welshfella):

No all you have to d is add 3 to both sides of 7 = x - 3 this will isolate x on right and give its value on left

OpenStudy (welshfella):

- weve done the -7x bit the next step is the +3.

OpenStudy (welshfella):

u still there?

OpenStudy (deathslasher99):

Yeah sorry

OpenStudy (deathslasher99):

So what was the answer?

OpenStudy (welshfella):

7 + 3 = x - 3 + 3 the rHS = x what is LHS?

OpenStudy (deathslasher99):

Im sorry i dont even know what that is lol

OpenStudy (deathslasher99):

I just wanted to solve this quick problem because of something i was calculating

OpenStudy (welshfella):

the reason we add 3 to both sides is because theres -3 on RHS and adding 3 will isolate x.

OpenStudy (welshfella):

surely you can work out what 7 + 3 is?

OpenStudy (deathslasher99):

ofcourse

OpenStudy (welshfella):

well there you go - thats the value of x.

OpenStudy (deathslasher99):

So then how would you write the answer as? Because i dont know where your going with this?

OpenStudy (deathslasher99):

What do you want to do with 10?

OpenStudy (deathslasher99):

Because remember there asking to solve it in (X,y) terms

OpenStudy (welshfella):

we have found x so now to find y plug the value of x into y = 7x + 7 and work it out

OpenStudy (deathslasher99):

Oh, i see

OpenStudy (welshfella):

y = 7(10) + 7

OpenStudy (welshfella):

the answer will be written as (10,y) where y is the result of the above

OpenStudy (deathslasher99):

So is it written as (10,77)?

OpenStudy (welshfella):

yea

OpenStudy (deathslasher99):

oh

OpenStudy (deathslasher99):

Ok

OpenStudy (welshfella):

x value always comes first

OpenStudy (deathslasher99):

Thank you, and before you go can i ask you something really quick?